An infinite line carries a linear uniform charge density of 19.6 nC/m). Determine the magnitude of the electric field ( in units of kN/C) at a distance of 5.0 cm from the line.


1
Expert's answer
2021-06-03T09:16:31-0400
"E=\frac{2\lambda k}{r}\\E=\frac{2(19.6\cdot10^{-9})(8.99\cdot10^{9})}{0.05}=7.0\frac{kN}{C}"